# From South to North: Leveraging Ground-Based LATs for Full-Sky CMB Delensing and Constraints on $r$

### v1

@@ 4.3 Results on the $r$ constraint
[…]
@@
Figure 11: The covariance matrix of the data vector $\mathbf{\hat{X}}_{b}$ for the LATS+SAT and LATN+LATS+SAT cases. Notice: The covariance values are shown in logarithmic scale for clarity. The nine slices represent the covariance matrices from bin 1 to bin 9, corresponding to the multipole range from $\ell=20$ to $\ell=200$, which are used for parameter constraints. We find an overall reduction in the matrix value for the LATN+LATS+SAT case (right panel).
@@
The summarized results from the Gradient-order method and Inverse-lensing method are presented in Table 5 and Table 6, respectively. For the LATS+SAT case, the uncertainty in $r$ decreases from approximately $0.72\times 10^{-3}$ to $0.57\times 10^{-3}$ after incorporating the LT as an independent observation channel into the likelihood. This corresponds to a reduction of about 21%. Furthermore, we find that incorporating LATN data leads to a further reduction in the uncertainty of $r$. Even without delensing, this inclusion yields nearly an 8% improvement. Remarkably, when the lensing template is constructed with LATN data included, the uncertainty in $r$ is reduced by an additional 19%—a level of improvement comparable to that achieved by incorporating external LSS tracers ($\sim$ 17%). Once LATN is included, the marginal gain from adding external data results in a further reduction of approximately 10% in the uncertainty. We also find a strong consistency between the two delensing methods for most cases, further reinforcing the robustness of our results.
@@
Table 5: The mean and $1\sigma$ standard deviation of each parameter using Gradient-order method, with the LATS+SAT case, LATS+SAT+External case, LATN+LATS+SAT case and LATN+LATS+SAT+External case. Gaussian approximation likelihood was used in likelihood analysis. Notice that in the main text, the terms ”Without delensing” and ”With delensing” are used interchangeably with ”Before adding LT” and ”After adding LT” shown here, respectively.
Gradient-order Method Parameter Input value Before adding LT (LATS+SAT) Before adding LT (LATN+LATS+SAT) After adding LT (LATS+SAT) After adding LT (LATS+SAT+External) After adding LT (LATN+LATS+SAT) After adding LT (LATN+LATS+SAT+External) $r(\times 10^{3})$ 0 0.002±0.7220.002$\pm$ 0.7220.002 ± 0.722 0.004±0.666-0.004$\pm$ 0.666- 0.004 ± 0.666 0.009±0.5680.009$\pm$ 0.5680.009 ± 0.568 0.002±0.4410.002$\pm$ 0.4410.002 ± 0.441 0.002±0.4270.002$\pm$ 0.4270.002 ± 0.427 0.002±0.3540.002$\pm$ 0.3540.002 ± 0.354 $A_{L}$ 1 1.000±0.0181.000$\pm$ $0$ 0.018 1.000±0.0171.000$\pm$ $0$ 0.017 1.000±0.0151.000$\pm$ $0$ 0.015 1.000±0.0131.000$\pm$ $0$ 0.013 1.000±0.0141.000$\pm$ $0$ 0.014 1.000±0.0121.000$\pm$ $0$ 0.012
Table 6: The mean and $1\sigma$ standard deviation of each parameter using Inverse-lensing method, with the LATS+SAT case and LATN+LATS+SAT case. Gaussian approximation likelihood was used in likelihood analysis. Notice that in the main text, the terms ”Without delensing” and ”With delensing” are used interchangeably with ”Before adding LT” and ”After adding LT” shown here, respectively. We do not present cases with +External since there is no reason to expect any evident difference in the Gradient-order method.
Inverse-lensing Method Parameter Input value Before adding LT (LATS+SAT) Before adding LT (LATN+LATS+SAT) After adding LT (LATS+SAT) After adding LT (LATN+LATS+SAT) $r(\times 10^{3})$ 0 0.002±0.7220.002$\pm$ 0.7220.002 ± 0.722 0.004±0.666-0.004$\pm$ 0.666- 0.004 ± 0.666 0.011±0.5710.011$\pm$ 0.5710.011 ± 0.571 0.007±0.4290.007$\pm$ 0.4290.007 ± 0.429 $A_{L}$ 1 1.000±0.0181.000$\pm$ $0$ 0.018 1.000±0.0171.000$\pm$ $0$ 0.017 1.000±0.0151.000$\pm$ $0$ 0.015 1.000±0.0141.000$\pm$ $0$ 0.014
4.4 Limitations of current forecasts
Although we have attempted to produce realistic forecasts to assess the improvement in constraining the tensor-to-scalar ratio $r$ from a ground-based large-aperture telescope in the Northern Hemisphere, one limitation of these forecasts is the bias introduced by non-Gaussian foregrounds. Such foreground residuals are known to bias internal lensing reconstruction by several percent [71, 72, 73, 74], due to their response to the quadratic estimator (QE). This bias propagates through the delensing procedure, ultimately affecting the inferred constraint on $r$ [56]. As the impact of these effects is expected to be sub-dominant for LAT surveys such as SO, we defer a detailed investigation to future work.
@@ 5 Conclusions
In this paper, we present a pipeline designed to improve full-sky constraints on $r$ through large-scale B-mode delensing for future ground-based and satellite CMB observations. As baseline, we consider a scenario (denoted as LATS+SAT), where a ground-based large-aperture telescope in the Southern Hemisphere (LATS) enables the detection of the lensing potential, while a satellite-based small-aperture telescope (SAT) facilitates the detection of large-scale polarization with a significantly low noise level. Furthermore, we find that incorporating an additional ground-based LAT in the Northern Hemisphere (denoted as LATN+LATS+SAT) provides two key advantages. On the one hand, the overlap between experiments helps reduce noise. On the other hand, the expanded sky coverage allows for a larger area where the lensing B-mode template can be constructed, thereby enhancing the effectiveness of B-mode delensing. These two factors together significantly improve the constraints on $r$, further reducing its uncertainty compared to the LATS+SAT case. We also incorporate the external LSS tracers to the delensing procedure as comparison.
We apply the NILC method to remove foreground components in the observations from both LATs and SAT, utilizing six frequency channels for LATs and fifteen for SAT. We observe a significant reduction in the power of the cleaned maps, owing to the method’s objective of minimizing the total variance. For polarization, we expect noise and foreground residuals to have minimal impact on the delensing procedure, as discussed in our prior work [75]. For temperature, it is well known that the ILC method is less effective in suppressing compact extragalactic foregrounds. In our analysis, we retain these residuals, as we expect them to have a negligible impact on our results. A more detailed investigation of this issue is left for future work.
@@
The lensing B-mode template (LT) is constructed using simulated data from the combined observations of the satellite mission (SAT) and ground-based LATs, employing two delensing approaches. For parameter estimation, we extend the likelihood function to incorporate all auto- and cross-bandpowers between the LT and the observed B-modes. Our results show that internal delensing using southern-hemisphere data alone (LATS+SAT) reduces the uncertainty in the tensor-to-scalar ratio $r$ by approximately 21% compared to the no-delensing case. Furthermore, the inclusion of northern LAT (LATN) enables full-sky internal delensing and yields an additional $\sim$19reduction in the uncertainty of $r$, comparable to the improvement from including external LSS tracers ($\sim$17%). Once LATN is incorporated, the marginal gain from LSS tracers diminishes to around 10%.

### v2

@@ 4.3 Results on the $r$ constraint
Table 4: Input value used for B-mode simulation and prior imposed on each parameters for MCMC sampling. $\mathcal{U}(a,b)$ denotes uniform distribution between $[a,b]$.
@@
Figure 11: The covariance matrix of the data vector $\mathbf{\hat{X}}_{b}$ for the LATS+SAT and LATN+LATS+SAT cases. Notice: Only the diagonal blocks ($\text{Cov}(C_{\text{bin}\_i^{th}}^{A},C_{\text{bin}\_i^{th}}^{B})$) are shown, and the covariance values are shown in logarithmic scale for clarity. The five slices represent the covariance matrices from bin 0 to bin 4, corresponding to the multipole range from $\ell=2$ to $\ell=200$, which are used for parameter constraints. We find an overall reduction in the matrix value for the LATN+LATS+SAT case (right panel).
@@
The summarized results from the Gradient-order method and Inverse-lensing method are presented in Table 5 and Table 6, respectively. For the LATS+SAT configuration, the uncertainty in the tensor-to-scalar ratio $r$ decreases from approximately $0.70\times 10^{-3}$ to $0.58\times 10^{-3}$ when the lensing template (LT) is incorporated as an independent observation channel in the likelihood, corresponding to an improvement of about 17%. We further find that the inclusion of LATN data leads to an additional reduction in the uncertainty of $r$. Even in the absence of delensing, adding LATN improves the constraint by nearly 11%. More remarkably, when the lensing template is constructed with LATN data included, the uncertainty in $r$ is reduced by a further 18%, from $0.58\times 10^{-3}$ to $0.45\times 10^{-3}$. This level of improvement is comparable to that achieved by incorporating external large-scale structure (LSS) tracers, which yields an approximately 13% reduction ($0.58\times 10^{-3}$ to $0.50\times 10^{-3}$). Once LATN data are included, the marginal benefit of adding external LSS information is correspondingly reduced, resulting in an additional improvement of about 10% in the uncertainty of $r$, from $0.45\times 10^{-3}$ to $0.38\times 10^{-3}$. We also find a strong consistency between the two delensing methods for most cases, further reinforcing the robustness of our results.
@@
Table 5: The mean and $1\sigma$ standard deviation of each parameter using Gradient-order method, with the LATS+SAT case, LATS+SAT+External case, LATN+LATS+SAT case and LATN+LATS+SAT+External case. HL likelihood and ”diag full” covariance conditioning were used in likelihood analysis. Notice that in the main text, the terms ”Without delensing” and ”With delensing” are used interchangeably with ”Before adding LT” and ”After adding LT” shown here, respectively.
Gradient-order Method Parameter Input value Before adding LT (LATS+SAT) Before adding LT (LATN+LATS+SAT) After adding LT (LATS+SAT) After adding LT (LATS+SAT+External) After adding LT (LATN+LATS+SAT) After adding LT (LATN+LATS+SAT+External) $r(\times 10^{3})$ 0 $021\pm 0.699 0.0$$25\pm 0.618 0.0$$14\pm 0.582 0.0$$25\pm 0.496 0.0$$16\pm 0.450 0.0$$031\pm 0.380$ $A_{L}$ 1 $1.000\pm 0.022$ $0\pm 0.020 1.001±0.0$$1\pm 0.017$ $1\pm 0.014 1.001±0.0$$1\pm 0.015$ $1.001\pm 0.012$
Table 6: The mean and $1\sigma$ standard deviation of each parameter using Inverse-lensing method, with the LATS+SAT case and LATN+LATS+SAT case. HL likelihood and ”diag full” covariance conditioning were used in likelihood analysis. Notice that in the main text, the terms ”Without delensing” and ”With delensing” are used interchangeably with ”Before adding LT” and ”After adding LT” shown here, respectively. We do not present cases with +External since there is no reason to expect any evident difference in the Gradient-order method.
Inverse-lensing Method Parameter Input value Before adding LT (LATS+SAT) Before adding LT (LATN+LATS+SAT) After adding LT (LATS+SAT) After adding LT (LATN+LATS+SAT) $r(\times 10^{3})$ 0 $021\pm 0.699 0.0$$25\pm 0.618 0.0$$06\pm 0.586 0.0$$003\pm 0.460$ $A_{L}$ 1 $1.000\pm 0.022$ $0\pm 0.020 1.001±0.0$$1\pm 0.016$ $1.001\pm 0.015$
@@ 5 Conclusions
In this paper, we present a pipeline designed to improve full-sky constraints on $r$ through large-scale B-mode delensing for future ground-based and satellite CMB observations. As baseline, we consider a scenario (denoted as LATS+SAT), where a ground-based large-aperture telescope in the Southern Hemisphere (LATS) enables the detection of the lensing potential, while a satellite-based small-aperture telescope (SAT) facilitates the detection of large-scale polarization with a significantly low noise level. Furthermore, we find that incorporating an additional ground-based LAT in the Northern Hemisphere (denoted as LATN+LATS+SAT) provides two key advantages. On the one hand, the overlap between experiments helps reduce the noise level. On the other hand, the expanded sky coverage increases the area over which the lensing B-mode template can be constructed, thereby enhancing the overall effectiveness of B-mode delensing. These two factors together significantly improve the constraints on $r$, further reducing its uncertainty compared to the LATS+SAT case. We also incorporate the external LSS tracers to the delensing procedure as comparison.
We apply the NILC method to remove foreground components in the observations from both LATs and SAT, utilizing six frequency channels for LATs and fifteen for SAT. We observe a significant reduction in the power of the cleaned maps, owing to the method’s objective of minimizing the total variance. For polarization, we expect noise and foreground residuals to have minimal impact on the delensing procedure, as discussed in our prior work [30]. For temperature maps, it is well established that the ILC-based cleaning is less effective at suppressing compact foregrounds. Due to their non-Gaussianity and correlations with the underlying matter distribution, residual extragalactic foregrounds can bias the internally-reconstructed lensing signal at the several-percent level [98, 89, 97, 38]. Such bias can then propagate into the delensing procedure and subsequently affect the likelihood analysis [17]. A similar bias may also arise from the non-Gaussianity of Galactic foregrounds [4, 20]. However, [4] demonstrates that, for an SO-like experiment employing MV HILC, the bias induced by Galactic foregrounds remains safely below the reconstruction uncertainty. Consistent conclusions are reported in [20], where no significant additional bias is found after foreground cleaning, enabling an unbiased measurement of the tensor-to-scalar ratio after delensing. In addition, our own analysis incorporating explicit foreground-mitigation strategies (Appendix D) yields constraints on $r$ that agree well with our baseline results. Based on these findings, we conclude that the baseline analysis remains robust against residual foreground contamination. Nonetheless, a more systematic optimization and evaluation of foreground-induced reconstruction biases will be pursued in future work.
@@
The lensing B-mode template (LT) is constructed using simulated data from the combined observations of the satellite mission (SAT) and ground-based LATs, employing two delensing approaches. For parameter estimation, we extend the likelihood function to incorporate all auto- and cross-bandpowers between the LT and the observed B-modes. Our results show that internal delensing using southern-hemisphere data alone (LATS+SAT) reduces the uncertainty in the tensor-to-scalar ratio $r$ by approximately 17% compared to the no-delensing case. Furthermore, the inclusion of northern LAT (LATN) enables full-sky internal delensing and yields an additional $\sim$18% reduction in the uncertainty of $r$, comparable to the improvement from including external LSS tracers ($\sim$13%). Once LATN is incorporated, the marginal gain from LSS tracers diminishes to around 10%.